# Tuning without Peeking: Provable Privacy and Generalization Bounds for LLM Post-Training

### v1

@@ Abstract
Gradient-based optimization is the workhorse of deep learning, offering efficient and scalable training via backpropagation. However, its reliance on large volumes of labeled data raises privacy and security concerns—such as susceptibility to data poisoning attacks—and the risk of overfitting. In contrast, black box optimization methods, which treat the model as an opaque function, relying solely on function evaluations to guide optimization, offer a promising alternative in scenarios where data access is restricted, adversarial risks are high, or overfitting is a concern. However, black box methods also pose significant challenges, including poor scalability to high-dimensional parameter spaces, as prevalent in large language models (LLMs), and high computational costs due to reliance on numerous model evaluations. This paper introduces BBoxER, an evolutionary black-box method for LLM post-training that induces an information bottleneck via implicit compression of the training data. Leveraging the tractability of information flow, we provide strong theoretical bounds on generalization, differential privacy, susceptibility to data poisoning attacks, and robustness to extraction attacks. BBoxER operates on top of pre-trained LLMs, offering a lightweight and modular enhancement suitable for deployment in restricted or privacy-sensitive environments, in addition to non-vacuous generalization guarantees. In experiments with LLMs, we demonstrate empirically that Retrofitting methods are able to learn, showing how a few iterations of BBoxER improve performance and generalize well on a benchmark of reasoning datasets. This positions BBoxER as an attractive add-on on top of gradient-based optimization.
@@ 1 Introduction
“It is impossible to produce superior performance unless you do something different from the majority” - Sir John Templeton
@@
Large Language Models (LLMs) have revolutionized natural language processing, achieving strong performance across diverse tasks through training on massive datasets – often hundreds of billions of tokens – and large-scale architectures [1, 2]. Yet LLMs suffer from persistent weaknesses, foremost among them vulnerability to extraction attacks [3, 4, 5, 6, 7] where training data can be leaked [8, 9], raising potential privacy concerns. They are also vulnerable to data poisoning [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20], and lack meaningful generalization guarantees at scale. These challenges are amplified in settings where data is scarce, sensitive, or adversarially curated, underscoring the need for more robust and privacy-aware optimization methods.
@@
Black-Box Optimization (BBO) covers methods suited to settings where gradient information is unreliable or unavailable (e.g., non-differentiable or noisy objectives). These algorithms are naturally robust to local minima and non-convexity. Although around for decades, BBO has been recognized only recently in machine learning, becoming a core tool in AutoML [21], for hyperparameter tuning, algorithm selection [22], and neural architecture search (NAS) [23]. Beyond random and grid search, techniques like Evolutionary Algorithms (e.g., CMA-ES [24, 25, 26], Differential Evolution [27], Particle Swarm Optimization [28]) and Bayesian Optimization [29, 30, 31] offer diverse trade-offs in scalability, sample efficiency, and exploration-exploitation dynamics. BBO methods still face limitations: they scale poorly to high-dimensional parameter spaces and require many function evaluations, making them unsuitable for full model training (except for Deep Reinforcement Learning, due to its high parallelism [32]). To reconcile the strengths and weaknesses of both paradigms, we adopt a hybrid strategy: using gradient-based pretraining for scale, followed by BBO on a small, targeted subset of the model to tackle privacy, poisoning, overfitting.
@@
Post-training introduces its own risks: instruction tuning and generation can lead to exploitative behavior [33, 16, 34], overfitting may harm robustness and generalization [35, 36, 37], and data poisoning can occur during reinforcement learning from human feedback (RLHF) [38] or direct preference optimization (DPO) [39, 10, 12, 13, 14, 15, 40, 41, 42]. This setting is well-suited for retrofitting – a concept initially developed to refine word vectors post-hoc using semantic knowledge [43], which we generalize here to LLMs for arbitrary downstream tasks.
@@
We introduce BBoxER (Black-Box Evolutionary Retrofitting), a comparison-based black-box retrofitting method [44] applicable after pretraining, fine-tuning, or reinforcement learning loops such as GRPO [45]. Positioned at the outermost layer of the pipeline, BBoxER requires no gradient access and integrates seamlessly with existing black-box libraries and algorithms [46]. The resulting sequence of queries and updates forms an implicit compression trace, enabling the use of compression-based generalization theory [47, 48], currently the only route to non-vacuous generalization bounds for LLMs [49]. It allows us to derive strong generalization and differential privacy guarantees that depend not on the number of modified parameters but on the complexity of the optimization trajectory. As a result, BBoxER ensures privacy by design, robustness to data poisoning, robustness to extraction attacks, and provable, non-vacuous generalization bounds to control overfitting. It is uniquely positioned to provide improvements from aggregate feedback of billions of users anonymously, without accessing individual data.
Since theory does not ensure that retrofitting leads to meaningful improvements, we validate our approach on multibillion LLMs across multiple reasoning benchmarks in verifiable domains (Sec. 5). Despite query budgets of only a few hundred model evaluations, BBoxER yields consistent gains and improved generalization, supporting its practical value as a post-hoc adaptation method for LLMs. BBoxER thus emerges as an alternative to standard fine-tuning, particularly in privacy-sensitive scenarios. To summarize, our contributions are as follows:
@@
    $\bullet$ BBoxER: We introduce a general-purpose BBO retrofitting framework for LLMs and flesh out its compression properties induced by comparison-based black-box optimization (Sec. 3.2), enabling safe and modular adaptation of pretrained and post-trained models (Sec. 3).
    $\bullet$ Theoretical generalization guarantees: We derive non-vacuous generalization bounds for retrofitted models. Unlike VC dimension [50] or Rademacher complexity [51] our bounds depend only on the complexity of the optimization trace (Sec. 4), and allow training budgets to scale linearly with dataset size, making them practical for LLMs (Sec. 4.1).
    $\bullet$ Privacy and robustness: Our framework ensures privacy by design and provides formal guarantees against poisoning (Sec. 4.2), and extraction attacks (Sec. 4.3).
    $\bullet$ Empirical validation in the low-budget regime: We evaluate BBoxER on GSM8K [52] and other math benchmarks [53, 54, 55, 56] using Llama3.1-8B [57] and Qwen-2.5-3B [58] base and instruct models, in a regime where our generalization bounds are non-vacuous. Despite tight query budgets, we observe consistent test improvements and positive transfer to out-of-distribution tasks (Sec. 5).
@@ 2 Background and Related Work
An extensive exposition of related work is provided in App. A.
@@
Data Poisoning. Recent poisoning attacks demonstrate that corrupting as little as 1-5% of fine-tuning or preference data can cause persistent and targeted behaviors in LLMs [15, 14], even across quantized variants [62]. These attacks evade standard detection and remain active during inference. Likewise, even the standard strategy of training on public data and fine-tuning on private data is susceptible to backdoor injection attacks [34]. In particular, [34] demonstrated that enabling neurons to hold a gradient for a single input and “deactivate” them to prevent overwriting during subsequent training leads to the model becoming vulnerable to gradient inversion attacks. Our approach follows a black-box optimization strategy and, consequently, by design, is secure against such an attack. While theoretical bounds [63] exist for simpler models, robust guarantees for LLMs remain elusive.
@@
3 Methods: Retrofitting with BBO
Retrofitting [68, 69, 70] is routinely deployed in industry, in order to adapt old devices to new contexts or new data. In machine learning, the term has mainly been used in natural language processing since Faruqui et al. [71], who adapted the representation of the word vector to take semantic knowledge into account. It is performed after classical training, i.e., without retraining the entire network. Videau et al. [44] adapted retrofitting to the black-box optimization context. In the present paper, we instantiate it as BBoxER in the specific case of Large Language Models. We consider comparison-based black-box optimization algorithms to obtain the explicit compression properties that lend themselves to the strong bounds proved in Sec. 4.
3.1 The Retrofitting Framework: BBoxER
Notation. Let ${\cal D}$ be the class of datasets of size $s$ drawn from a probability distribution $F\in{\cal F}$ over a domain ${\cal X}\times{\cal Y}$. Models map $X$ to $Y$. Let $\omega$ be the random seed corresponding to the possibly randomized optimization algorithm and $\omega_{D}$ the random seed associated with the draw of $D\in{\cal D}$.
Algorithm 1 BBoxER($m_{0},\omega,D,a,b$): Returns a modification of initial model $m_{0}$ by BBO algorithm $a$ (see Alg. 2) on dataset $D$ with budget $b$ and random seed $\omega$.
1:  $I_{0}=a.Initialize(\omega)$
@@ 3.1 The Retrofitting Framework: BBoxER
9:  $FinalModel=modified(m_{0},\widehat{x}=a.recommend())$
Initial model and modifiers.
We denote the initial model by $m_{0}$, irrespective of its creation. In the context of LLMs, $m_{0}$ is a checkpoint of publicly available LLMs such as [57]; refer to [44] for examples of other initial deep learning models. $modified(m_{0},x)$ refers to a modification of $m_{0}$ parametrized by some $x$. We prove all results for an arbitrary $modified$ function. For example, $modified$ can refer to a multiplicative low-rank update of a particular attention layer of the LLM; $x$ then contains the two low-rank matrices that parametrize the update (see App. B).
BBoxER. Alg. 1 presents our abstract retrofitting framework. It features, as a key component, $a$, a black-box optimization algorithm. We use the ask-and-tell formalism [72] and object-oriented notation to keep track of the internal state of $a$ and make the link to compression: $a$ is an object specified by its methods $ask$, $tell$, $recommend$, $compare$, and $numCases$.
A few observations are in order: First, Algorithm $a$ is initialized and returns its internal state $I_{0}$ in 1. Then at each iteration, Algorithm $a$, via $a.ask$ at 3, proposes $x_{i}$ for next modification. 4 introduces the branching factor $k_{i}$ of Algorithm $a$ through $a.numCases$. $k_{i}$ is the finite number of possible outputs of $a.compare$, that we assume to be upper bounded. At 5, $a.compare$ returns $choice_{i}\in\{1,\dots,k_{i}\}$, the result of comparing various previously proposed models $m_{1},\dots,m_{i}$ on the data $D$. For example, for a simple algorithm based on comparing $m_{i}$ with the best model so far $m_{best}$, $compare$ returns $1$ if $m_{i}$ is better than $m_{best}$ and 00 otherwise. In 6, $a$ is updated via $a.tell$, the only method that modifies $a$. Thus, the history of arguments to $a.tell$ allows to recover the entire internal state $I_{i}$ of $a$. It is entirely determined by $\omega,a,b,choice_{1},\dots,choice_{i}$, updated at 7, and contains all the information available to the optimization algorithm. More precisely, the number of distinct possible outputs of BBoxER is upper bounded by the number of distinct possible values of $I_{b}$. Finally $a.recommend$ at 9 returns the modification $\widehat{x}$ to obtain the final model. It depends on $\omega,a$ (its internal state), $b,choice_{1},\dots,choice_{b}$.
We assume that algorithm $a$ is deterministic given the seed $\omega$; stochastic algorithms can be obtained by randomizing the seed $\omega$ in $a$.
BBoxER: Link with compression. In  Alg. 1, $FinalModel=modified(m_{0},\hat{x})$ is a deterministic function of $m_{0},\omega,a,b,D$. More precisely, by induction on $b$, $\widehat{x}$ only depends on $\omega,a,b,choice_{1},\dots,choice_{b}$. Therefore, we can construct $\widehat{x}$ without $D$, and there exists a mapping X
@@
But $(choice_{1},\dots,choice_{b})$ is a deterministic function of $m_{0},\omega,a,b,D$. Hence, there exists a deterministic $compression$ function $c$ defined by
@@
Here, $X$ mimics the BBoxER Alg. 1 but returns $\widehat{x}$ instead of $modified(m_{0},\widehat{x})$, and uses $c$ instead of $D$. The dataset, conditioned on $c=(choice_{1},\dots,choice_{b})$, is entirely removed from Equation 1. To flesh out BBoxER as compression, we can merge Equation 1 and Equation 2 into:
@@
$FinalModel$ only depends on the data through the compression bottleneck $c$ of bounded size (see  Figure 2, Appendix, for illustration). The impact of this compression is formalized in Sec. 4, and used to prove guarantees in terms of overfitting, robustness w.r.t privacy and robustness.
3.2 BBoxER: The link with comparison-based black-box optimization
BBO is concerned with optimizing objective functions for which the only available information is the value they take at points of the search space. A first motivation for using BBO is to be able to use non-differentiable loss functions for post-training, such as signals from A/B testing. In the BBoxER context, BBO also only requires inferences of the LLM at hand, lightening the computational cost of post-training. But the main motivation of this work is to benefit from the compression brought by BBoxER (Equation 3): this adds the constraint of a finite branching factor (number of possible answers of the optimization algorithm, 4 of Alg. 1), a cornerstone of all theoretical results of Sec. 4.
This constraint de facto eliminates most, if not all (see Sec. D.2), derivative-free optimization algorithms from Mathematical Programming [73], but also Bayesian Optimization [29, 31]. Most remaining BBO algorithms can be seen as flavors of Evolutionary Algorithms (EAs) [74], the most prominent being Evolution Strategies (ESs) [75], Differential Evolution (DE) [27] and Particle Swarm Optimization (PSO) [28], to name a few (more in Sec. D.1). They are all stochastic population-based algorithms, they differ in the way they explore the search space but they are comparison-based: after choosing new points to evaluate in the search space, they only need the results of comparison with previously sampled points, and don’t make use of the actual values of the objective function– this is key for a finite branching factor needed in Sec. 4.
All experiments presented in this work have been run on the Open Source platform for BBO Nevergrad [46], and are described in Sec. D.1,while App. E details the derivation of their average branching factors, summarized in Tab. 3. Let us briefly survey the most popular BBO algorithms. (1+1)-EAs (like OneFifth, Lengler, Discrete used in the experiments Sec. 5) evolve one point, that is randomly modified, and the best of initial and modified points is kept, giving an obvious branching factor of 2 (if ties are broken, e.g., in favor of the new point). DE and PSO evolve a full population, but each new point is generally compared to the points it was moved from – again a branching factor of 2 (or 3 for the Curr-to-best variants). Covariance Matrix Adaptation (CMA-ES) [25, 76] and its D-CMA variant evolve a full population of size $\mu$ using a linear recombination and an adaptive Gaussian mutation (hence the name) and their branching factor is either 2 or $2\sqrt{\mu}$ depending on the variant (different weights for the linear recombination).
Nevergrad also includes Bet-and-Run algorithms, that run in parallel N algorithms with finite branching factor, each with budget $b/N$: the global branching factor remains finite and can even be smaller than that of the base algorithms (Sec. E.4). For instance, for $N=3$ and if all branching factors are less than $2$ the average branching factor of the Bet-and-Run combination is ($b]{\left(3\cdot 2^{b/3}\right)}\ll 2$.

### v2

@@ Abstract
Gradient-based optimization is the workhorse of deep learning, offering efficient and scalable training via backpropagation. However, exposing gradients during training can leak sensitive information about the underlying data, raising privacy and security concerns such as susceptibility to data poisoning attacks. In contrast, black box optimization methods, which treat the model as an opaque function, relying solely on function evaluations to guide optimization, offer a promising alternative in scenarios where data access is restricted, adversarial risks are high, or overfitting is a concern. This paper introduces BBoxER, an evolutionary black-box method for LLM post-training that induces an information bottleneck via implicit compression of the training data. Leveraging the tractability of information flow, we provide non-vacuous generalization bounds and strong theoretical guarantees for differential privacy, robustness to data poisoning attacks, and extraction attacks. In experiments with LLMs, we demonstrate empirically that black-box optimization methods—despite the scalability and computational challenges inherent to black-box approaches—are able to learn, showing how a few iterations of BBoxER improve performance, generalize well on a benchmark of reasoning datasets, and are robust to membership inference attacks. This positions BBoxER as an attractive add-on on top of gradient-based optimization, offering suitability for deployment in restricted or privacy-sensitive environments while also providing non-vacuous generalization guarantees.
@@ 1 Introduction
Large Language Models (LLMs) have revolutionized natural language processing, achieving strong performance across diverse tasks through training on massive datasets – often hundreds of billions of tokens – and large-scale architectures [1, 2]. Yet LLMs suffer from persistent weaknesses, foremost among them vulnerability to extraction attacks [3, 4, 5] where training data can be leaked [6, 7], raising potential privacy concerns. They are also vulnerable to data poisoning [8, 9], and lack meaningful generalization guarantees at scale. These challenges are amplified in settings where data is scarce, sensitive, or adversarially curated, underscoring the need for more robust and privacy-aware optimization methods.
@@
Black-Box Optimization (BBO) covers methods suited to settings where gradient information is unreliable or unavailable (e.g., non-differentiable or noisy objectives). These algorithms are naturally robust to local minima and non-convexity. Although around for decades, BBO has been recognized only recently in machine learning, becoming a core tool in AutoML [10], for hyperparameter tuning, algorithm selection [11], and neural architecture search (NAS) [12]. Beyond random and grid search, techniques like Evolutionary Algorithms (e.g., CMA-ES [13, 14, 15], Differential Evolution [16], Particle Swarm Optimization [17]) and Bayesian Optimization [18, 19, 20] offer diverse trade-offs in scalability, sample efficiency, and exploration-exploitation dynamics. BBO methods still face limitations: they scale poorly to high-dimensional parameter spaces and require many function evaluations, making them unsuitable for full model training (except for Deep Reinforcement Learning, due to its high parallelism [21]). To reconcile the strengths and weaknesses of both paradigms, we adopt a hybrid strategy: using gradient-based pretraining for scale, followed by BBO on a small, targeted subset of the model to tackle privacy, poisoning, and overfitting.
@@
Post-training introduces its own risks: instruction tuning and generation can lead to exploitative behavior [22, 23], overfitting may harm robustness and generalization [24], and data poisoning can occur during reinforcement learning from human feedback (RLHF) [25, 26] or direct preference optimization (DPO) [27, 28]. This setting is well-suited for retrofitting – a concept initially developed to refine word vectors post-hoc using semantic knowledge [29], which we generalize here to LLMs for arbitrary downstream tasks.
@@
We introduce BBoxER (Black-Box Evolutionary Retrofitting), a comparison-based black-box retrofitting method [30] applicable after pretraining, fine-tuning, or reinforcement learning loops such as GRPO [31]. Positioned in the outermost layer of the pipeline, BBoxER requires no gradient access and integrates seamlessly with existing black-box libraries and algorithms [32]. The resulting sequence of queries and updates forms an implicit compression trace, enabling the use of compression-based generalization theory [33, 34], currently the only route to non-vacuous generalization bounds for LLMs [35]. It allows us to derive strong generalization and differential privacy guarantees that - unlike VC dimension [36] or Rademacher complexity [37] - do not depend on the number of modified parameters but on the complexity of the optimization trajectory. As a result, BBoxER ensures privacy by design, robustness to data poisoning, robustness to extraction attacks, and provable, non-vacuous generalization bounds to control overfitting. It is uniquely positioned to provide improvements from aggregate feedback of billions of users anonymously, without accessing individual data.
While our approach establishes strong theoretical guarantees, guarantees alone would be meaningless if they came at the cost of utility—after all, doing nothing also ensures perfect privacy. What is striking with BBoxER is that, even under tight query budgets of only a few hundred model evaluations, it consistently improves performance and generalization on billion-scale LLMs across reasoning benchmarks. In other words, the framework delivers provable generalization and safety without sacrificing usefulness, turning theory into a practical tool for post-hoc adaptation of LLMs, especially in privacy-sensitive settings.
Our contributions. BBoxER introduces a new perspective on LLM post-training: instead of fine-tuning weights or relying on reinforcement learning, we show that simple black-box retrofitting can achieve safe and modular adaptation. Building on this idea, we make three main contributions:
@@
$\bullet$ A general-purpose retrofitting framework. We formalize BBoxER (Alg.˜1), a comparison-based black-box optimization scheme that compresses optimization traces, enabling safe and modular adaptation of pretrained and post-trained LLMs (Sec.˜3).
$\bullet$ Strong theoretical guarantees. We derive non-vacuous generalization bounds (Corollary˜1) that scale linearly with dataset size, and establish privacy by design through provable robustness to poisoning (Equation˜7) and resistance to extraction attacks (Corollary˜2) (Secs.˜4, 4.2 and 4.3).
$\bullet$ Empirical validation on billion-scale LLMs. Despite tight query budgets, we show consistent gains on GSM8K and related math benchmarks with Llama3.1-8B and Qwen-2.5-3B, and provide empirical evidence supporting our theoretical claims: retrofitted models with BBoxER resist Membership Inference Attacks, unlike fine-tuned counterparts at equal utility (Sec.˜5).
@@ 2 Background and Related Work
An extensive exposition of related work is provided in App.˜A.
@@
Data Poisoning. Recent poisoning attacks show that corrupting as little as 1–5% of fine-tuning or preference data can induce persistent, targeted behaviors in LLMs [26, 25], even in quantized variants [41]. These attacks evade standard detection and persist during inference. Moreover, the common strategy of training on public data then fine-tuning on private data is also vulnerable to backdoor injection [42]. Specifically, Feng and Tramèr [42] showed that allowing neurons to retain a gradient from a single input and later “deactivate” to avoid overwriting makes models susceptible to gradient inversion. Our black-box optimization approach is, by design, secure against such attacks. While theoretical bounds exist for simpler models [43], robust guarantees for LLMs remain elusive.
Privacy and Robustness. LLMs risk leaking private or confidential content [4]. Differential privacy (DP) techniques [44] like DP-SGD [45, 46] and PATE [47] are the standard paradigm for privacy preserving learning, offering formal guarantees but suffering from utility and scalability challenges, due to the necessary per-sample gradient clipping and adding noise that is inversely proportional to $\varepsilon$. Our method avoids the need for noisy gradient updates by enforcing a compression bottleneck through comparison-based optimization. For instance, in the specific case of ensuring the privacy of user input prompts and outputs in A/B testing, it achieves $(\varepsilon=0,\delta=0)$.
3 Methods: Retrofitting with BBO
Retrofitting [48, 49, 50] is routinely deployed in industry, in order to adapt old devices to new contexts or new data. In machine learning, the term has mainly been used in natural language processing since Faruqui et al. [51], who adapted the representation of the word vector to take semantic knowledge into account. It is performed after classical training, i.e., without retraining the entire network. Videau et al. [30] adapted retrofitting to the black-box optimization context. In the present paper, we instantiate retrofitting as BBoxER and consider comparison-based black-box optimization algorithms to obtain the explicit compression properties that lend themselves to the strong bounds proved in Sec.˜4.
A key motivation for using BBO is that it allows post-training with non-differentiable loss functions such as accuracy or signals from A/B testing. And since it only requires model inference, it significantly reduces memory usage. More importantly, the main motivation of this work is to leverage the compression achieved by reducing the entire dataset’s forward pass to just a single—or a few—comparison results, yielding a finite branching factor. This constrains our deployment of BBO to comparison-based algorithms, which include Evolutionary Algorithms (EA) and Strategies (ES) like Population-Based Evolution Strategies (e.g., CMA, DiagonalCMA), Differential Evolution, and Particle Swarm Optimization - all described in Sec.˜D.2. While a vast class of BBO algorithms is not comparison based (see Sec.˜D.2 for an anthology of BBO) this nonetheless gives us a substantial variety of BBO algorithms to deploy. For example, the (1+1)-OneFifth strategy samples new points from a Gaussian centered on the current best solution and adjusts the mutation step size based on the acceptance rate. Similarly, CMA-ES updates both the mean and covariance of its sampling distribution based on comparisons within the current population. More algorithms are described in App.˜D. Our BBoxER framework is modular and algorithm agnostic, and we instantiate it with methods from the Nevergrad suite [32].
3.1 The Retrofitting Framework: BBoxER
Notation. Let $\cal D$} be the class of datasets of size $s$ over a domain ${\cal X}\times{\cal Y}$. Models map $\cal X$} to $\cal Y$}. Let $\omega$ be the random seed corresponding to the possibly randomized optimization algorithm $a$ and $\omega_{D}$ the random seed associated with the draw of $D\in{\cal D}$.
Initial model and modifiers.
We denote $m_{0}$ the initial model, irrespective of its creation. $modified(m_{0},x)$ refers to a modification of $m_{0}$ parametrized by some $x$. We prove all results for an arbitrary modifiedmodified function. For example, modifiedmodified can refer to a multiplicative low-rank update of a particular attention layer of the LLM, $x$ then contains the two low-rank matrices that
parametrize the update (see App.˜B).
Alg.˜1 presents our abstract retrofitting framework BBoxER.
Algorithm 1 BBoxER($m_{0},\omega,D,a,b$): Returns a modification of initial model $m_{0}$ by BBO algorithm $a$ on dataset $D$ with budget $b$ and random seed $\omega$.
@@ 3.1 The Retrofitting Framework: BBoxER
[…]
First, Algorithm $a$ is initialized and returns its initial internal state $I_{0}$ in ˜1. Then for $b$ iterations, Algorithm $a$, via $a.ask$ at ˜3, proposes a new sample $x_{i}$, which used to obtain a new model $m_{i}$ through modifiedmodified. ˜4 introduces the branching factor $k_{i}$, representing the finite number of possible outputs of $a.compare$, which we assume is upper bounded. At ˜5, $a.compare$ returns $choice_{i}\in\{1,\dots,k_{i}\}$, the outcome of comparing various previously proposed models $m_{1},\dots,m_{i}$ on the data $D$. For instance, the comparison may involve the current model and the best model obtained so far, which corresponds to $k_{i}=2$. The outcome of this comparison can be based on user-provided preference, such as in A/B testing, or on performance evaluations over $D$, where we compute a metric (e.g., accuracy) but only retain the binary comparison result (e.g., 0 or 1). In ˜7, internal state of $a$ is updated to account for the comparison result. Finally $a.recommend$ at ˜9 returns the modification $\widehat{x}$ used to obtain the final model.
It is important to note that the internal state $I_{b}$ of algorithm $a$ is entirely determined by $\omega,a,b,$ $choice_{1},\dots,choice_{b}$ and contains all the information needed to obtain $\widehat{x}$, and thus the final modified model. Moreover, the number of distinct possible outputs of BBoxER is upper bounded by the number of distinct possible values of $I_{b}$. We assume that the randomness of algorithm $a$ is fixed given the seed $\omega$; stochastic algorithms can be obtained by randomizing the seed $\omega$ in $a$.
3.2 BBoxER: Link with compression
[…]
@@ 3.2 BBoxER: Link with compression
But $(choice_{1},\dots,choice_{b})$ is a deterministic function of $m_{0},\omega,a,b,D$. Hence, there exists a deterministic compressioncompression function $c$ defined by
@@
[…]
@@
FinalModelFinalModel only depends on the data through the compression bottleneck $c$ of bounded size (see Figure˜2, Appendix, for illustration). The impact of this compression is formalized in Sec.˜4, and used to prove guarantees in terms of overfitting, privacy robustness w.r.t poisoning and extraction attacks.